5.60 g of an unknown gas at STP fills a 500. mL flask. What is the molar mass of the gas?
Answer #1

From ideal gas equation we know, PV = nRT

All the terms have their usual significance.

PV =w/M * RT ( where , n is number of moles = w/M, we = weight of gas, M = molecular weight of the gas)

Therefore, PV = wRT/M

Or, M = wRT/PV . . . . . . . (i)

At standard conditions , value of pressure, P = 1atm and standard temperature, T = 273K

And value of universal gas constant, R = 0.082 lit. atm. K-1.mol-1

Given that, weight of gas,w = 5.60g and volume of gas, V = 500mL = 0.5 litre

Therefore, substituting all the value, we get

M = 5.60 * 0.082 * 273 /1 * 0.5

= 250.7232g/mole